This is a practical appliaction of quadratic equations.
How would I set up the following solution to this problem?

The hudson river flows at a rate of 3 miles per hour. A patrol boat travels 60 miles upriver, and returns in a total time of 9 hours. What is the speed of the boat in still water?


I realy don't know how to thank the tutor's that helped me on the last question.I am new to this. So thank you to the tutors that helped me last time. I would like to say thank you in advance to the tutor that will help me on this question.Your help is appreciated.

Answer by scott8148(6628)

let x="speed in still water" __ d=rt __ t=d/r __ time upstream plus time downstream equals 9 hr

upstream __ t=60/(r-3)

downstream __ t=60/(r+3)

[60/(r-3)]+[60/(r+3)]=9 __ multiplying by (r+3)(r-3) __ 60(r+3)+60(r-3)=9(r+3)(r-3)

120r=9r^2-81 __ 0=9r^2-120r-81 __ dividing by 3 __ 0=3r^2-40r-27